# Ultrafast and Ultralight Network-Based Intelligent System for Real-time Diagnosis of Ear diseases in Any Devices

### 21 Jan 2024 v2

@@ Abstract
Traditional ear disease diagnosis heavily depends on experienced specialists and specialized equipment, frequently resulting in misdiagnoses, treatment delays, and financial burdens for some patients. Utilizing deep learning models for efficient ear disease diagnosis has proven effective and affordable. However, existing research overlooked model inference speed and parameter size required for deployment. To tackle these challenges, we constructed a large-scale dataset comprising eight ear disease categories and normal ear canal samples from two hospitals. Inspired by ShuffleNetV2, we developed Best-EarNet, an ultrafast and ultralight network enabling real-time ear disease diagnosis. Best-EarNet incorporates the novel Local-Global Spatial Feature Fusion Module which can capture global and local spatial information simultaneously and guide the network to focus on crucial regions within feature maps at various levels, mitigating low accuracy issues. Moreover, our network uses multiple auxiliary classification heads for efficient parameter optimization. With 0.77M parameters, Best-EarNet achieves an average frames per second of 80 on CPU. Employing transfer learning and five-fold cross-validation with 22,581 images from Hospital-1, the model achieves an impressive 95.23% accuracy. External testing on 1,652 images from Hospital-2 validates its performance, yielding 92.14% accuracy. Compared to state-of-the-art networks, Best-EarNet establishes a new state-of-the-art (SOTA) in practical applications. Most importantly, we developed an intelligent diagnosis system called Ear Keeper, which can be deployed on common electronic devices. By manipulating a compact electronic otoscope, users can perform comprehensive scanning and diagnosis of the ear canal using real-time video. This study provides a novel paradigm for ear endoscopy and other medical endoscopic image recognition applications.
Keywords Ear Disease  $\cdot$ Deep Learning  $\cdot$ Feature Fusion  $\cdot$ Light-weight Network  $\cdot$ Intelligent Application
@@ 3.2.2 Local-Global Spatial Feature Fusion Module
3.2.3 Processing of branch path and Multiple auxiliary classification heads
For Flow, it is not directly passed into LGSFF but processed through a separate branch path. Firstly, an Average Pooling Layer and a Point-wise Convolution Layer are applied to resize the feature map of Flow from 48×28×48 to 384×7×7. Then, DropBlock is used to perform regularization on the feature map. After that, Efficient Channel Attention (ECA)[31] is applied to enable the network to focus more on the low-level feature information in Flow. Batch Normalization and ReLU activation functions are applied to the Flow features. The internal structure of ECA is shown in the diagram, where K=5. Additionally, we add multiple auxiliary classification heads to Best-EarNet. Compared to a single loss, using multiple auxiliary classification heads introduces additional losses from the shallow layers of the network, which increases the gradient signals propagated during backpropagation. This provides extra regularization, improves model robustness, and helps the network converge better and optimize its parameters. The internal structure of these auxiliary heads is shown in the Fig. 2d.
@@ 5 Conclusion
Acknowledgments
This paper is supported by the NSF of Guangdong Province (No.2022A1515011044, No.2023A1515010885), and the project of promoting research capabilities for key constructed disciplines in Guangdong Province (No.2021ZDJS028).

### 10 Apr 2024 v4

# Ear-Keeper: Real-time diagnosis of ear lesions utilizing ultralight-ultrafast ConvNet and large-scale ear endoscopic dataset

@@ Abstract
Deep learning-based ear disease diagnosis technology has proven effective and affordable. However, due to the lack of ear endoscope datasets with diversity, the practical potential of the deep learning model has not been thoroughly studied. Moreover, existing research failed to achieve a good trade-off between model inference speed and parameter size, rendering models inapplicable in real-world settings. To address these challenges, we constructed the first large-scale ear endoscopic dataset comprising eight types of ear diseases and disease-free samples from two institutions. Inspired by ShuffleNetV2, we proposed Best-EarNet, an ultrafast and ultralight network enabling real-time ear disease diagnosis. Best-EarNet incorporates a novel Local-Global Spatial Feature Fusion Module and multi-scale supervision strategy, which facilitates the model focusing on global-local information within feature maps at various levels. Utilizing transfer learning, the accuracy of Best-EarNet with only 0.77M parameters achieves 95.23% (internal 22,581 images) and 92.14% (external 1,652 images), respectively. In particular, it achieves an average frame per second of 80 on the CPU. From the perspective of model practicality, the proposed Best-EarNet is superior to state-of-the-art backbone models in ear lesion detection tasks. Most importantly, Ear-keeper, an intelligent diagnosis system based Best-EarNet, was developed successfully and deployed on common electronic devices (smartphone, tablet computer and personal computer). In the future, Ear-Keeper has the potential to assist the public and healthcare providers in performing comprehensive scanning and diagnosis of the ear canal in real-time video, thereby promptly detecting ear lesions.
Keywords Ear Diseases  $\cdot$ Deep Learning  $\cdot$ Feature Fusion  $\cdot$ Light-weight Network  $\cdot$ Intelligent System
@@ 3.2.2 Local-Global Spatial Feature Fusion Module
3.2.3 Processing of branch path and multi-scale supervision
For Flow, it is not directly passed into LGSFF but processed through a separate branch path. Firstly, an Average Pooling Layer and a Point-wise Convolution Layer are applied to resize the feature map of Flow from 48×28×48 to 384×7×7. Then, DropBlock is used to perform regularization on the feature map. After that, Efficient Channel Attention (ECA)[31] is applied to enable the network to focus more on the low-level feature information in Flow. Batch Normalization and ReLU activation functions are applied to the Flow features. The internal structure of ECA is shown in the diagram, where K=5.
Multi-scale supervision is particularly important for processing complex medical images, as these images may contain structures and details on multiple scales. Therefore, to encourage the network to learn discriminative features at multiple levels, we use multiple auxiliary classification heads to supervise feature representations at different levels, which helps improve the learning efficiency and final performance of the entire network. Concretely, we add auxiliary classification heads to stage-2’s ouput and stage-3’s ouput of Best-EarNet. Compared to a single classification head, using multiple auxiliary classification heads introduces additional losses from the shallow layers of the network, which increases the gradient signals propagated during backpropagation. This provides extra regularization, improves model robustness, and helps the network converge better and optimize its parameters. The internal structure of these auxiliary heads is shown in the Fig. 2d.